# DiffEyeSyn: Diffusion-based User-specific Eye Movement Synthesis

### v1

@@ 6. Discussion
6.1. Importance of User-Specific Eye Movement Synthesis and DiffEyeSyn Applications
Previous research has primarily focused on modelling human eye movement behaviour at low frequencies, typically below 30 Hz (Jiao et al., 2024; Wang et al., 2023), often overlooking the user-specific information embedded in high-frequency components. In contrast, we propose the first computational approach to modelling user-specific eye movement behaviour. User-specific eye movement synthesis is a promising research field due to its wide-ranging application potential.
@@ 6.1. Importance of User-Specific Eye Movement Synthesis and DiffEyeSyn Applications
Furthermore, user-specific information plays an important role in tasks such as gaze data imputation and gaze data super-resolution. These tasks address challenges such as missing data due to pupil detection fails or limited sampling frequencies of eye tracking sensors. While interpolation methods have achieved good time-series metrics, such as mean absolute error, in filling missing data, they lack the ability to provide user-specific solutions (Jiao et al., 2023). DiffEyeSyn directly addresses these gaps. By incorporating observed eye movement sequences and target user embeddings, DiffEyeSyn complements interpolation methods by offering personalised, high-frequency eye movement data. For example, in Section 5.2 and 5.3, we demonstrated DiffEyeSyn’s ability in an extreme user-specific gaze data super-resolution and imputation task (upsampling 20 Hz eye movements to 1,000 Hz). DiffEyeSyn synthesised eye movements achieve high cosine similarities in both tasks. However, we acknowledge that current outcomes (Figure 5) show that while DiffEyeSyn effectively captures user-specific information of the target user, the synthesised eye movements exhibit noticeable spatial variations from the input data. For tasks like user-specific gaze data super-resolution, the user-specific information and the spatial gaze directions are equally important. Therefore, we did not compare DiffEyeSyn with gaze data super-resolution methods (Jiao et al., 2023) and other traditional interpolation approaches that focus on generating spatially-controlled eye movements ignoring user-specific information. This is due to most gaze-based user authentication models working on the velocity domain, with a powerful user authenticator working on the spatial domain, DiffEyeSyn can be easily updated to an efficient gaze data super-resolution and imputation method.
Beyond these fundamental applications, DiffEyeSyn’s capabilities extend to areas like character animation and gaze-based context recognition. For instance, DiffEyeSyn facilitates high-frame-rate animation of natural eye movements while also enabling the customisation of eye movements for distinct characters.
6.2. The performance upper bound of DiffEyeSyn
@@ 6.3. Evaluating Synthetic Eye Movements
Since there are no established metrics in the eye tracking domain to evaluate the realism of synthetic eye movements, following prior work (Prasse et al., 2023), we evaluated the synthetic velocity distributions by calculating $JS$ divergence. Besides, we provided qualitative results in Section 5 and supplementary materials to visually evaluate the realism of the synthetic data. Moreover, we introduce to use pretrained user authenticators to evaluate the performance in synthesising user-specific information. Instead of using different pretrained models for both training and evaluating our method, we used the EKYT model in both stages. Because these pretrained models are deep learning models, different models for the same task might not share common thoughts for the same input. Using the state-of-the-art pretrained model for both training and evaluations is common in the computer vision community (Kim et al., 2022, 2023; Yang et al., 2023; Li et al., 2023). For instance, in DiffusionCLIP (Kim et al., 2022), authors used the CLIP model in their training loss and also evaluated the CLIP score. However, we acknowledge that the model trained in this way might be overfitting to the limitations of the selected pretrained model. A better user authentication model with less limitations can also make the evaluation more precise.
@@ 6.4. Limitation and Future Work
We evaluated our method on the GazeBase (Griffith et al., 2021) and JuDo1000 (Makowski et al., 2020) datasets. These datasets have similar quality characteristics, such as being collected using an Eyelink stationary eye tracker at a very high frequency (1,000 Hz) with participants’ heads fixed. This similarity allows for a robust evaluation of synthetic high-frequency user-specific eye movements. In recent years, eye tracking-enabled head-mounted devices, such as Apple Vision Pro and Meta Quest, have become popular commercial products. Eye movement-based user authentication has significant potential to become a regular unlock method for such headsets. However, head-mounted devices typically employ low-frequency eye trackers, and the recorded eye movements include both head and eye movements. Using DiffEyeSyn to synthesise eye movements for training user authentication systems for head-mounted devices requires additional efforts. This includes retraining the user authenticator on a virtual reality (VR) or augmented reality (AR) dataset and subsequently using this retrained user authenticator to finetune DiffEyeSyn. We will investigate this in future work.
@@ 7. Conclusion
This paper presented DiffEyeSyn, the first computational approach for user-specific eye movement synthesis. We conducted extensive experiments to evaluate the performance of DiffEyeSyn in user identity recovery and user identity manipulation, demonstrating DiffEyeSyn can synthesise realistic user-specific eye movement data that is indistinguishable from human eye movements. In particular, we also demonstrated that DiffEyeSyn can be directly used in user-specific gaze imputation and gaze data super-resolution. We believe that DiffEyeSyn sets the basic groundwork for user-specific eye movement synthesis that has significant application potential, such as for character animation, eye movement biometrics, or gaze-based activity and context recognition.

### v2

@@ 6. Discussion
6.1. Importance of User-Specific Eye Movement Synthesis and Applications
Previous research has primarily focused on modelling human eye movement behaviour at low frequencies, typically below 30 Hz (Jiao et al., 2024; Wang et al., 2023), often overlooking the user-specific subtle movements embedded in high-frequency eye tracking data. In contrast, we propose the first computational approach to modelling user-specific eye movement behaviour. User-specific eye movement synthesis is a promising research field due to its wide-ranging application potential.
@@ 6.1. Importance of User-Specific Eye Movement Synthesis and Applications
Furthermore, user-specific information plays an important role in tasks such as gaze data imputation. While interpolation methods have achieved good time-series metrics, such as mean absolute error, in filling missing data, they lack the ability to provide user-specific solutions (Jiao et al., 2023). DiffEyeSyn directly addresses these gaps. By incorporating observed eye movement sequences and target user embeddings, DiffEyeSyn complements interpolation methods by offering personalised, high-frequency eye movement data. For example, in Section 5.2 and 5.3, we demonstrated DiffEyeSyn’s ability in an extreme user-specific gaze data super-resolution and imputation task (upsampling 20 Hz eye movements to 1,000 Hz). DiffEyeSyn synthesised eye movements achieve high cosine similarities in both tasks. However, we acknowledge that current outcomes (Figure 6) show that while DiffEyeSyn effectively captures user-specific information of the target user, the synthesised eye movements may exhibit minor spatial variations from the input data, depending on the target eye movement sequence. While this is acceptable for tasks like user-specific eye movement synthesis, where subtle user-specific eye movements are key, it may limit performance in spatially-sensitive applications. Since many authentication models rely on velocity features rather than spatial ones, we did not directly compare DiffEyeSyn with interpolation methods or super-resolution techniques that only optimise for spatial accuracy.
Beyond these core applications, DiffEyeSyn also supports tasks like personalised eye movement animation. For example, it can be used to animate virtual characters with natural, user-specific gaze behaviour at high frame rates (see Figure 1 and supplementary video).
6.2. Performance Upper Bound
@@ 6.3. Evaluating Synthetic Eye Movements
Traditional scanpath metrics are not suitable for evaluating our task. Prior work has shown that such metrics perform poorly in assessing synthetic eye movements and fail to align with expert visual judgement (Jiao et al., 2024). Moreover, scanpath prediction differs fundamentally from our objective. While scanpath models aim to predict fixation locations and durations for a given visual stimulus, DiffEyeSyn focuses on generating subtle, individual-specific gaze patterns independent from visual stimulus. A key advantage of DiffEyeSyn is its ability to inject these subtle movements into any existing eye movement sequence, including scanpaths. As such, these tasks are complementary: DiffEyeSyn can serve as a post-processing module to personalise the outputs of scanpath prediction models.
Table 2 shows that adding high-frequency components from the ground truth to an identity-removed signal improves performance but still falls short of DiffEyeSyn in recovering user-specific information. This highlights that user identity is not tied solely to event types or frequency components, but rather embedded in the fine-grained dynamics of gaze behaviour—patterns likely shaped by each individual’s physiological eye system. As user-specific information is encoded in these subtle variations, existing quantitative metrics are insufficient for evaluation. At present, the most reliable approach is to assess downstream tasks where user-specificity is critical. Gaze-based user authentication and identification are prime examples of such tasks. Accordingly, we propose using a pretrained user authenticator to evaluate the ability of DiffEyeSyn to preserve identity-related information and apply DiffEyeSyn in a downstream task of augmenting gaze-based user identification.
We adopt the EKYT model (Lohr and Komogortsev, 2022), a state-of-the-art method for gaze-based user authentication, both for training and evaluation. While using the same model across both stages risks overfitting to its particular biases, this practice is common in the vision community. For example, DiffusionCLIP (Kim et al., 2022) and other recent works (Kim et al., 2023; Yang et al., 2023; Li et al., 2023) use the CLIP model in both training objectives and evaluation metrics. However, we acknowledge that the model trained in this way might be overfitting to the limitations of the selected pretrained model. We anticipate that stronger and more general gaze encoders will emerge in future, allowing for more robust evaluation of user-specific synthesis methods.
In addition to downstream performance, we follow prior work (Prasse et al., 2023, 2024) and report Jensen–Shannon ($JS$) divergence between the velocity distributions of synthetic and real eye movements. We also provide qualitative results in Section 5 and in the supplementary materials, enabling visual inspection of realism. Our overall goal is to simulate realistic human eye movements. DiffEyeSyn represents an important first step towards this goal and demonstrates state-of-the-art performance in capturing velocity distributions—an aspect that is largely overlooked by existing scanpath methods.
@@ 6.4. Limitation and Future Work
We evaluated our method on the GazeBase (Griffith et al., 2021) and JuDo1000 (Makowski et al., 2020) datasets. These datasets have similar quality characteristics, such as being collected using an Eyelink stationary eye tracker at a very high frequency (1,000 Hz) with participants’ heads fixed. This similarity allows for a robust evaluation of synthetic user-specific eye movements. In recent years, eye tracking-enabled head-mounted devices, such as Apple Vision Pro and Meta Quest, have become popular commercial products. Eye movement-based user authentication has significant potential to become a regular unlock method for such headsets. However, head-mounted devices typically employ low-frequency eye trackers, and the recorded eye movements include both head and eye movements. Using DiffEyeSyn to synthesise eye movements for training user authentication systems for head-mounted devices requires additional efforts. This includes retraining the user authenticator on a virtual reality (VR) or augmented reality (AR) dataset and subsequently using this retrained user authenticator to fine-tune DiffEyeSyn. We will investigate this in future work.
@@ 7. Conclusion
This paper presented DiffEyeSyn, the first computational approach for user-specific eye movement synthesis. We conducted extensive experiments to evaluate the performance of DiffEyeSyn in user identity recovery and user identity manipulation, demonstrating DiffEyeSyn can synthesise realistic user-specific eye movement data that contains user-specific characteristics. In particular, we also demonstrated that DiffEyeSyn can be directly used in gaze-based user identification. We believe that DiffEyeSyn sets the basic groundwork for user-specific eye movement synthesis that has significant application potential, such as for personalised character animation, eye movement biometrics, and user-specific gaze imputation.